# Cell Library Characterization for Composite Current Source Models Based on Gaussian Process Regression and Active Learning

### v1

@@ Abstract.
copyright: nonejournalyear: 2024copyright: acmlicensedconference: 2024 ACM/IEEE International Symposium on Machine Learning for CAD; September 9–11, 2024; Salt Lake City, UT, USAbooktitle: 2024 ACM/IEEE International Symposium on Machine Learning for CAD (MLCAD ’24), September 9–11, 2024, Salt Lake City, UT, USAdoi: 10.1145/3670474.3685965isbn: 979-8-4007-0699-8/24/09footnotetext: This work was supported in part by the National Natural Science Foundation of China under Grant 62174031 and in part by the Fundamental Research Funds for the Central Universities.footnotetext: Both authors contributed equally to this research.footnotetext: Peng Cao is the corresponding author.
@@ 2.1. Related Works
The deep neural network (DNN) model is applied to perform standard cell characterization for nominal timing (8965191, ) and statistical timing (naswali2021fast, ), respectively, which use data from sparse characterization to generate delay models at required sign-off corners but ignore the relation with transistor-level cell topology. A characterization approach using ridge regression for aging-aware cell library is detailed in (9394564, ), which enables designers to assess their circuits under precisely selected degradation scenarios. Similarly, the work of (8988626, ) introduced an innovative logical circuit simulation method implemented using neural networks based on the current source model. Furthermore, the research in (chen2023unleashing, ) employs a multi-layer perceptron (MLP) model to create power supply noise (PSN) waveform for each standard cell. Although various ML algorithms have been utilized, the method for selecting training data of cell characterization has not been thoroughly studied. This oversight has a significant impact on prediction accuracy and the simulation cost required to generate the data. Moreover, the uncertainty of predicted results by ML algorithms may degrade the accuracy of cell library, which is vital for circuit design.

### v2

@@ Abstract.
copyright: nonejournalyear: 2024doi: 10.1145/3670474.3685965isbn: 979-8-4007-0699-8/24/09footnotetext: Both authors contributed equally to this research.footnotetext: Peng Cao is the corresponding author.
@@ 2.1. Related Works
The deep neural network (DNN) model is applied to perform standard cell characterization for nominal timing (8965191, ) and statistical timing (naswali2021fast, ), respectively, which use data from sparse characterization to generate delay models at required sign-off corners but ignore the relation with transistor-level cell topology. A characterization approach using ridge regression for aging-aware cell library is detailed in (9394564, ), which enables designers to assess their circuits under precisely selected degradation scenarios. Similarly, the work of (8988626, ) introduced an innovative logical circuit simulation method implemented using neural networks based on the current source model. The research in (chen2023unleashing, ) employs a multi-layer perceptron (MLP) model to create power supply noise (PSN) waveform for each standard cell. Furthermore, (zhou2025lvfgen, ) utilizes GPR with AL to estimate statistical timing distributions in OCV characterization.
Although various ML algorithms have been utilized, the method for selecting training data of CCS characterization has not been thoroughly studied. This oversight has a significant impact on prediction accuracy and the simulation cost required to generate the data. Moreover, the uncertainty of predicted results by ML algorithms may degrade the accuracy of cell library, which is vital for circuit design.